OpenStudy (anonymous):

Simplify. square root of 6 (square root of 33 + 7)

OpenStudy (anonymous):

Is this your question:\[\sqrt{6}(\sqrt{33}+7)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Just distribute root(6) to the brackets..

OpenStudy (anonymous):

when two terms are raised to the same root you are able to multiply whatever is under the square root: \[\sqrt{a}\sqrt{b}=\sqrt{ab}\]

OpenStudy (anonymous):

So, this now becomes \[\sqrt{6*33}+7\sqrt{6}\]\[\sqrt{198}+7\sqrt{6}\]

OpenStudy (anonymous):

\[\sqrt{22*9}+7\sqrt{6}\] \[\sqrt{22}*\sqrt{9}+7\sqrt{6}\] \[3\sqrt{22}+7\sqrt{6}\]

OpenStudy (anonymous):

Do you understand @tngirl865

OpenStudy (anonymous):

yes I do.. thank u very much for helping me!